OpenStudy (anonymous):

Solve for y. −3(8 − y) = 6(1/2 y – 4) A. y = –24 B. y = 8/3 C. infinitely many solutions D. no solution @iambatman @jagr2713

jagr2713 (jagr2713):

can you draw this please

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):



OpenStudy (isaiah.feynman):

You can start by dividing both sides by 3

jagr2713 (jagr2713):

Multiply −3 by each term inside the parentheses (8−y). −3(8)−3(−y)=3y−24 Multiply −3 by the 8 inside the parentheses. −3⋅8−3(−y)=3y−24 Multiply −3 by 8 to get −24. −24−3(−y)=3y−24 Multiply −3 by the −y inside the parentheses. −24−3⋅−y=3y−24 Multiply −3 by −y to get 3y. −24+3y=3y−24 Reorder the polynomial −24+3y alphabetically from left to right, starting with the highest order term. 3y−24=3y−24 3y−24=3y−24 Since 3y contains the variable to solve for, move it to the left-hand side of the equation by subtracting 3y from both sides. 3y−24−3y=−24 Since 3y and −3y are like terms, add −3y to 3y to get 0. 0−24=−24 Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. −24=−24 Since −24=−24, am i close?? SO MUCH TYPING

jagr2713 (jagr2713):

@Isaiah.Feynman HELP

OpenStudy (isaiah.feynman):

Divide both sides by -3.

jagr2713 (jagr2713):

oh was i close?

OpenStudy (isaiah.feynman):